# Make the Fastest Faster: Importance Mask for Interactive Volume Visualization using Reconstruction Neural Networks

### v1

@@ 2.2 Differentiable Compaction and Decompaction
The goal of our work is to derive an optimal $S_{im}$ from $S_{sp}$, however, the shape of the sampling pattern varies for different RecNN. For example, super-resolution RecNN uses a downsampling pattern while the foveated rendering RecNN uses a foveal pattern. The downsampling pattern is evenly distributed across the PR-Image. The foveal pattern generally refers to the region of an image where the eye’s fovea (central vision) is focused, which is where human vision is most detailed. In foveated rendering, only the pixels in this high-resolution foveal area are rendered in full detail, while the surrounding areas corresponding to peripheral vision are rendered at a lower resolution. To uniformly process sampling patterns with different shapes, we introduce a compaction layer to compact the pixels in the sampling pattern into a regular 2D matrix. Fig. 2 show how the PR-Image is selectively rendered according to downsampling and foveal patterns, and how the PR-Image is compacted into a compact 2D matrix. We name this 2d matrix a compact image (C-Image). The compaction collects each pixel within the sampling pattern by following a straightforward Raster scan order on the PR-Image. If the C-Image is not filled up after compaction, zeros are padded to its empty pixels. We can see that the C-Image of super-resolution RecNN is the corresponding low-res version of the FR-Image, while the C-Image of foveated rendering RecNN is more irregular. The compaction layer is a practice of dataset distillation where only informative pixels are gathered for efficient training. Decompaction is the reverse of the compaction. The compaction and decompaction layers are 2D point transformations between PR-Image space and C-Image space. To make the compaction and decompaction layers differentiable, transformation is implemented as a translation matrix multiplying each pixel in one space to map its value to the transformed location of another space. The benefit of incorporating compaction and decompaction layers includes: 1) Reduce the dimension of the redundant training data into a compact subset. 2) Decrease the number of parameters of the IML network. 3) Improve the speed of convergence during training. 4) Improve the speed of inferencing for interactive visualization. 5) Allow the proposed IML and IMS Networks to remain agnostic to diverse sampling patterns within a unified network architecture.
Figure 2: Compaction and decompation for super-resolutin RecNN and foveated rendering RecNN.
@@ 2.3 Importance Mask Learning Network
The functions of IML Net include: 1) For each rendered sampling pattern, learn an IM that selects only the important pixels from it to render for the downstream reconstructions. 2) Learn a decoder that reconstructs the full rendering of the sampling pattern (C-Image) from the partially rendered sampling pattern (PR-C-Image) based on the learned IM.
Figure 3: Network architecture of IML Net. IML Net works on a compact image (C-Image) domain which has a resolution of $n\times n$. The full-res output has a resolution of $m\times m$ where $m>n$.
@@ 2.3.1 Network Architecture
Fig. 3 shows the network architecture of IML Net. IML Net is an autoencoder neural network that takes a C-Image after compaction as input and outputs an image as close as the C-Image. The number of pixels and their distribution on C-Image are defined by the specific sampling pattern selected. The importance mask is learned on top of it by optimizing the network parameters through backpropagation during training. Key components of the network are:
Encoder and Decoder: The first and the last subnetworks of the IML Net are encoder and decoder sharing the same network structure as shown in Fig. 4. We utilize standard U-Net [27] to construct the encoder/decoder network. Our IML Net is invariant to the input C-image resolution due to the use of convolutional layers. For the encoder, C_in is set to 3 for taking RGB C-Image after compaction as input, C_out is set to 1 for outputting a grayscale image where the value of each pixel correlates to the probability of being an important pixel. The U-Net structure of the encoder can effectively learn such pixel values to find the IM. After the last layer of the encoder, a softplus layer is used to convert the output value into the positive range for the following normalization layer. For the decoder, both C_in and C_out are set to 3 for taking the partially rendered C-Image (PR-C-Image) as input and outputting an image close to the input C-Image of the encoder. The decoder of our IML network performs a low-level in-painting to complete the missing pixels in the PR-C-Image to reconstruct the complete C-Image.
Normalization: The output of the encoder presents preliminary information about the importance of each pixel, and we name this output Preliminary Importance Mask (PIM). We apply a similar practice proposed by Weiss et al. [44] to normalize the PIM with a hyperparameter, called mean value, which will control how much percentage of C-Image pixels as important pixels. It is worth noticing that our method is trying to solve the problem of interactively generating visualization for novel views from rendered IM while their work [44] provides a differential sampling method that helps the learning of IM. The normalized PIM, can be computed:
@@
where the $i$ and $j$ are pixel indices, $u$ is the prescribed mean value and $l$ is the lower bound on the sample distribution which is set to 0. $u_{PIM}$ is the mean of PIM. $\delta$ is a small constant set as $71e-$ 7 to avoid zero division. All the above operations ensure that the normalization layer remains differentiable. It is worth noticing that while the normalization step can also be performed using a softmax layer, this approach may result in a loss of control over the proportion of C-Image pixels designated as important through the prescribed mean. Our normalization step enables the hyperparameter tuning to investigate the relationship between the sparsity of important pixels and reconstruction performance.
Rejection Sampling: The rejection sampling layer is to generate a binary IM from the continuous PIM’. The practice is filtering out the important pixel through a comparison between the pixel value in PIM’ and a randomly sampled value. The rationale behind this practice is due to spatial correlation found in the prediction from image generative neural networks. This fact makes the direct learning of the important pixel as a blob of connected pixels which doesn’t distinguish the underlying structure. Comparing the prediction with an uncorrelated random pattern can solve this issue and create isolated pixels that reflect the importance of the underlying structure. The random pattern $P$ is a grayscale image with the same dimension as C-Image with its pixel value randomly sampled from a uniform distribution between 0 and 1. We use plastic sampling to generate $P$ for the superior quality of the learned IM compared to other sampling strategies[44]. The IM can be retrieved through an independent Bernoulli process through rejection sampling.
@@
$IM_{ij}=Sigmoid(\alpha(PIM^{\prime}_{ij}-P_{ij})),\;Sigmoid(x)=\frac{1}{1+e^{-x}}$ (6)
where $\alpha$ is a positive scalar determining the steepness of the Sigmoid function. The pixel value of each IM is in the range $[0,1]$. The above equation is only used for training the IML Net. When using IML Net during the inferencing, which will be discussed in Section 2.4.4, only the regular rejection sampling of Eq. 5 is used. In that case, a binary mask using 1 bit per pixel is enough to represent the importance mask.
@@
During the inferencing stage, this layer is replaced by actual pixel-level rendering computation defined by the chosen volume rendering algorithm.
@@ 2.3.2 Loss Functions
The IML Net can be trained by itself with a compaction loss function which is an image loss function between the input C-Image and reconstructed C-Image. This practice is suitable for optimizing the off-the-shelf pre-trained RecNNs without going through the long end-to-end training process resulting from adding a complex downstream RecNN. The IML Net can also be trained together with downstream RecNN selected for various sampling patterns. In this case, the loss function is the weighted sum of the previous compacting loss and the image loss between the predicted reconstructed full-resolution image (FR-Image) and the ground truth (GT) image rendered using the baseline rendering algorithm without using RecNN. This end-to-end training will give better reconstruction results by jointly optimizing the parameters from both the IML Net and the downstream RecNN. In our work, we use the Mean Squared Error (MSE) as the image loss function and 0.5 as the weight for both image losses if trained end-to-end.
@@ 2.4.1 Network Architecture
The focus of this paper is on visualizing scientific data, which is a time-sensitive application. Therefore, we need to keep the main steps of the pipeline, including training and inferencing, finished as fast as possible. Our training includes two steps: training the IML Net and training the IMS Net. We provide a method to speed up the training of IML network, as mentioned in Section 3.6.4, by training a standalone IML Net without connecting to the downstream RecNN. However, we cannot carry out the same optimization on training the IMS Net. As a result, we have to select a GAN with less training/inferencing complexity. DCGAN [25] provides a better trade-off between the complexity and the reconstruction quality than other GAN architectures like cGAN (Conditional GAN) [14] and StyleGAN [18] whose training times are too long for scientific visualization. Fig. 5 shows the network architecture of IMS Net and the discriminator used for training. IMS Net is a generator that takes a sparse set of view parameters to generate a 2D binary image as the IM. The IMS Net is trained through the generative adversarial network (GAN). Key components of the network are:
@@
$view=(eye\_x/y/z,\;lookAt\_x/y/z,\;up\_x/y/z)$ (8)
where the $eye$ is the camera position, $lookAt$ is the focal point where the camera is looking, and $up$ is the "up" direction of the camera. IMS Net is a DCGAN with its detail structure shown in Fig. 6(a). It consists of several transposed convolutional layers with batch normalization and ReLU activation functions.
Discriminator: Fig. 6(b) shows the detailed structure of the binary classifier as a discriminator. The discriminator consisted of several convolutional layers with batch normalization and Leaky ReLU activation functions.
@@ 2.4.2 Loss Functions
As shown in Fig. 5, we consider two losses to train the IMS Net: image loss on the IM and adversarial loss. The adversarial loss is composed of two parts: the generator loss and the discriminator loss. The discriminator loss (cross-entropy of the binary classifier) measures how well the discriminator can distinguish between real and generated images.
$L_{D}=-\mathbb{E}{x\sim p{data}(x)}[\log D(x)]-\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (9)
where $D(\mathbf{x})$ is the discriminator’s output for a real image. $D(G(\mathbf{z}))$ is the discriminator’s output for a fake image $G(\mathbf{z})$. The generator loss measures how well the generator can fool the discriminator.
@@
The loss function selection and configuration are crucial in training a GAN. Detail results using different loss functions for training IMS Net are discussed in Section 3.5.2.
2.4.3 Training
  • Avoid learning the IM directly from the sparse view parameters but from C-image which presents more structural information. Our experiments show the IM directly learned from view parameters is a binary mask whose important pixels are evenly distributed across the region of the volumetric object. In that case, important pixels couldn’t properly capture the locations of informative regions of the rendered image.
  • Make the generator learn to predict simple patterns. The IM that the IMS Net tries to learn is a simple 2D binary image with only a single channel and finite outputs (-1 or 1).
  • Scale the input binary IM from $\{0,1\}$ to $\{-1,1\}$.
  • Use hyperbolic tangent function (Tanh) at the end of the generator to constrain the output between -1 and 1.
  • Use adversarial loss together with advanced perceptual loss [16] instead of simple BCE or MSE loss. Perceptual loss is a loss function that measures the difference between high-level feature representations of images rather than their pixel-wise difference. It leverages pre-trained deep neural networks (e.g., VGG [32]) to compare images at a feature level, focusing on perceptual quality and structural similarity. Perceptual loss evaluates the similarity between two images based on the semantic content, such as textures, edges, and patterns, offering superior perceptual quality compared to pixel-wise loss like MSE and MAE.
2.4.4 Inferencing
Once the IMS Net is trained, a visualization image synthesis network is constructed by combining three components: the trained IMS Net, the decoder of the trained IML Net, and the RecNN as shown in Fig. 7. We name this network IML/S Net + RecNN. The new network directly generates visualization image from the new view with improved rendering latency than the original RecNN. Detailed quality and timing evaluation of IML/S Net + RecNN is discussed Section 3.6.
Figure 7: Network architecture of the proposed visualization image synthesis network, IML/S Net + RecNN. IML/S is the IMS Net connected with decoder of IML Net. Trained IML/S Net will be used for image synthesis to directly predict FR-Image from view parameters.
3 Experiments and Evaluation
3.1 Datasets
3.1.1 Volumetric Datasets
For quality and performance evaluation, we select 4 large-scale volume datasets with distinct spatial features collected from diverse domains. The Chameleon dataset is CT scan of a chameleon. The Beechnut dataset is a microCT scan of a dried beechnut. The Rayleigh-Taylor dataset is a time step of a density field in a simulation of the mixing transition in Rayleigh-Taylor instability. The Flame dataset is a simulated combustion 3D scalar field. subsubsection 3.1.2 shows the detailed information of the testing datasets. The spacings of all datasets are normalized within the spatial range [-1, 1] with values normalized within the range [0, 1].
3.1.2 Training Data
The training dataset used to train the networks is the view and fully rendered image (FR-Image) pairs. First, we randomly select 2000 views around the volume of each volumetric dataset. Second, a FR-Image is rendered using specific volume rendering algorithm with lighting effect. In this work we select commonly used ray casting DVR as the renderer. Third, filter the FR-Image with the sampling pattern to generate the partially rendered image (PR-Image) where only pixels in the sampling pattern are kept. Fourth, apply compaction on PR-Image to generate C-Image. For training the IML Net standalone, the input is C-Image while the output is the predicted C-Image from decoder. For training the IML Net end-to-end with the RecNN, the input is the C-Image while the output is the predicted FR-image from RecNN. For training the IMS Net, the input is the view parameters while the output is the learned IM derived from the trained IML Net. The training and validation partition ratio is 9:1. Each volumetric dataset also generates a testing dataset consisting of 120 views forming an exploratory trajectory to simulate real user exploration. The FR-Image is rendered with a resolution of $512\times 512$ using the Visualization Toolkit (VTK) [30] with a customized time stamp to measure the rendering time for each pixel. The interpolation is set to trilinear. The sample distance is 0.02. The CPU thread is set as a single thread to faithfully measure the time duration when rendering individual pixel in sequence.
Figure 8: IM prediction quality comparison using different configurations of loss functions on Chameleon dataset.
@@ 3.1.2 Training Data
{tabu} c c c c Dataset Resolution Size Data Type
Chameleon $1024^{2}\times 1088$ 2.1 GB uint16
Beechnut $1024^{2}\times 1546$ 3.0 GB uint16
Rayleigh-Taylor $1024^{3}$ 4.0 GB float32
Flame $1408\times 1080\times 1100$ 6.23 GB float32
3.2 Experimental Setup
The experiments are designed to investigate the interactive volume visualization using RecNN. We are going to evaluate both the quality and performance using traditional volume visualization using RecNN and the proposed image synthesis network, IML/S Net + RecNN. The input is a sequence of unseen views and the output are the FR-images. We use PSNR as the main metric to evaluate rendering quality. The computing platform is a desktop featuring an Intel(R) Core(TM) i7-7700K CPU with 8 threads running at 4.20GHz, paired with 16GB of DDR4 DRAM clocked at 3200MHz, and operating on Ubuntu 20.04.4 LTS.
3.3 RecNN Selection and Configuration
We select two state-of-the-art RecNNs used for interactive volume visualization, EnhanceNet and FoVolNet, to showcase how our method can help to further improve their rendering latencies. For the super-resolution RecNN using EnhanceNet, we also use 4×4$\times$4 × super-resolution setting, the same as the experiment in its paper. Since our evaluation only focuses on one-shot image generation, for fair comparison, we remove the recurrent connections of the FoVolNet and only keep its W-Net structure for static image prediction. Fixed foveated rendering, where the high-resolution region is static and fixed in the center of the screen, is used in the experiment rather than dynamic foveated rendering. The foveal pattern $FP$ is generated by the following methods:
where $P_{ij}$ random pattern generated by blue noise which doesn’t generate unevenly distributed samples from low-frequency energy spikes in the spatial domain[3]. The $E_{ij}$ is a modulation surface to control the size of the foveated area through $\sigma$:
$E_{ij}=e^{-0.5(f_{x}^{2}+f_{y}^{2})\sigma}$ (12)
where $f_{x}$ and $f_{y}$ are the center coordinate of the foveal pattern.
(a) IMS Net
(b) IML/S Net + EnhaneNet
Figure 9: Qualitative comparison of IM prediction and reconstruction using different configurations of loss functions on Chameleon dataset.
3.4 Training
The proposed neural networks are trained using the PyTorch software stack to accelerate the training and inferencing performance on a single NVIDIA RTX A6000 GPU. The Adaptive Moment Estimation (Adam) optimizer is used with an initial learning rate of 0.001 for training both IML and IMS Nets. A validation dataset is selected to detect the overfitting on the training dataset and terminate the network parameters from updating through the early stop mechanism. The patience of the early stop is set as 20 epochs for IML Net and 1000 for IMS Net. The IML Net can be trained standalone or trained end-to-end with RecNN.
3.5 Ablation Study
3.5.1 Network Configuration
Encoder and Decoder of IML: The encoder of the IML Net is critical to learning an accurate IM for it provides a comprehensive guess of the IM. Its U-Net architecture is capable of distinguishing fine-grained details through its skip connections and recommending candidates of important pixels. Compared with the encoder that serves as an information filter, the decoder serves as an information reconstruction that reconstructs the PR-C-Image to a full C-Image. We investigate how the network configuration of the IML encoder and decoder affects the performance of the proposed rendering pipeline. Since the decoder of the IML Net mirrors the encoder’s configuration, we modify both together by adjusting the depth of their U-Nets. We fixed the architecture of IMS Net and only adjusted the IML Net. As the depth of the encoder and decoder increases, its reconstruction quality increases sublinearly (as shown in Fig. 9(b) and Fig. 9(c)) while the number of parameters of the network increases superlinearly (as shown in Fig. 9(a)). While a deeper IML Net can provide better reconstruction quality, its larger size slows down inference, leading to higher rendering latency. We select the depth of both the encoder and decoder as 5 for its balanced performance between rendering quality and inferencing latency.
@@ 3.5.1 Network Configuration
Generator of IMS: The generator of IMS Net takes a novel view and generates the corresponding IM. The accuracy of the generated IM determines the quality of the input to the downstream IML decoder for reconstruction. We evaluate the reconstruction capability of IMS Net by using different numbers of transposed convolutional layers. We fixed the architecture of IML Net and only adjusted the IMS Net. As the number of transposed convolutional layers increases, similarly to adjusting the depth of IML Net, its reconstruction quality increases sublinearly (as shown in Fig. 10(b) and Fig. 10(c)) while the number of parameters of the network increases superlinearly (as shown in Fig. 10(a)). We also observed that adjusting the number of transposed convolutional layers in IMS Net has a greater impact on the reconstruction quality than the depth of the U-Net in IML Net. We select the layer size as 8 for IMS Net to balance rendering quality and inferencing latency.

### v2

@@ 2.2 Differentiable Compaction and Decompaction
The goal of our work is to derive an optimal $S_{im}$ from $S_{sp}$, however, the shape of the sampling pattern varies for different RecNN. For example, super-resolution RecNN uses a downsampling pattern while the foveated rendering RecNN uses a foveal pattern. The downsampling pattern is evenly distributed across the PR-Image. The foveal pattern generally refers to the region of an image where the eye’s fovea (central vision) is focused, which is where human vision is most detailed. In foveated rendering, only the pixels in this high-resolution foveal area are rendered in full detail, while the surrounding areas corresponding to peripheral vision are rendered at a much lower resolution. To uniformly process sampling patterns with different distributions, we introduce a compaction layer to compact the pixels in the sampling pattern into a regular 2D matrix. Fig.˜2 shows how the PR-Image is selectively rendered according to downsampling and foveal patterns, and how the PR-Image is compacted into a regular 2D matrix. We name this 2d matrix a compact image (C-Image). The compaction collects each pixel within the sampling pattern by following a straightforward Raster scan order on the PR-Image. If the C-Image is not filled after compaction, zeros are padded to its empty pixels. We can see that the C-Image of super-resolution RecNN is the corresponding low-res version of the FR-Image, while the C-Image of foveated rendering RecNN is more irregular. The compaction layer is a practice of dataset distillation where only informative pixels are gathered for efficient training. Decompaction is the reverse of compaction. The compaction and decompaction layers are 2D point transformations between PR-Image space and C-Image space. To make the compaction and decompaction layers differentiable, a transformation is implemented as a translation matrix multiplying each pixel in one space to map its value to the transformed location of another space. During compaction, the C-Image is calculated through the linear transformation:
$C{\text{-}}Image=PR{\text{-}}Image\times C_{RecNN}$ (4)
Where $C_{RecNN}$ is the Compaction Matrix, which is a 2D constant matrix defined by the specific RecNN. During decompaction, the PR-Image can be retrieved:
$PR{\text{-}}Image=C{\text{-}}Image\times D_{RecNN},\quad D_{RecNN}=C_{RecNN}^{-1}$ (5)
Where the Decompaction Matrix $D_{RecNN}$ is the inverse of the $C_{RecNN}$. The benefit of incorporating compaction and decompaction layers includes: 1) Reduce the dimension of the redundant training data into a compact subset for efficient Meta Learning. 2) Decrease the number of parameters of the IML network. 3) Improve the speed of convergence during training. 4) Improve the speed of inferencing for interactive visualization. 5) Allow the proposed IML and IMS Networks to remain agnostic to diverse sampling patterns within a unified network architecture. Compared to the LAS, the proposed compaction/decompaction enables learning from a low-res intermediate embedding (C-Image), therefore, further reducing the expensive pixel calculation for faster rendering.
@@ 2.3 Importance Mask Learning Network
The functions of IML Net include: 1) For each rendered sampling pattern, learn an IM that selects only the important pixels from it to render for the downstream reconstructions. 2) Learn a decoder that reconstructs the full rendering of the C-Image from the partially rendered C-Image (PR-C-Image) based on the learned IM.
@@ 2.3.1 Network Architecture
Fig.˜3 shows the network architecture of IML Net. IML Net is an autoencoder neural network that takes a C-Image after compaction as input and outputs an image as close as the C-Image. The number of pixels and their distribution on C-Image are defined by the specific sampling pattern selected. The IM is learned on top of it by optimizing the network parameters through backpropagation during training. Key components of the network are:
Encoder and Decoder: The first and the last subnetworks of the IML Net are encoder and decoder sharing the same network structure as shown in Fig.˜4. We utilize a standard U-Net [27] to construct the encoder/decoder network. Our IML Net is invariant to the input C-image resolution due to the use of convolutional layers. For the encoder, C_in is set to 3 for taking the RGB C-Image after compaction as input, C_out is set to 1 for outputting a grayscale image where the value of each pixel correlates to the probability of being an important pixel. The U-Net structure of the encoder can effectively learn such pixel values to find the IM. After the last layer of the encoder, a softplus layer is used to convert the output value into the positive range for the following normalization layer. For the decoder, both C_in and C_out are set to 3 for taking the partially rendered C-Image (PR-C-Image) as input and outputting an image close to the input C-Image of the encoder. The decoder of our IML network performs a low-level in-painting to complete the missing pixels in the PR-C-Image to reconstruct the complete C-Image.
Normalization: The output of the encoder presents preliminary information about the importance of each pixel, and we name this output Preliminary Importance Mask (PIM). We apply a similar practice proposed by Weiss et al. [45] to normalize the PIM with a hyperparameter, called mean value, which will control the percentage of C-Image pixels selected as important pixels. The normalized PIM, can be computed:
@@
where the $i$ and $j$ are pixel indices, $u$ is the prescribed mean value, and $l$ is the lower bound on the sample distribution, which is set to 0 in this work. $u_{PIM}$ is the mean of PIM. $\delta$ is a small constant set as $10^{-7}$ to avoid zero division. All the above operations ensure that the normalization layer remains differentiable. It is worth noticing that while the normalization step can also be performed using a softmax layer, this approach may result in a loss of control over the proportion of C-Image pixels designated as important through the prescribed mean. Our normalization step enables the hyperparameter tuning to investigate the relationship between the sparsity of important pixels and reconstruction performance.
Rejection Sampling: The rejection sampling layer is to generate a binary IM from the $PIM^{\prime}$ with continuous values. The practice is filtering out the important pixel through a comparison between the pixel value in $PIM^{\prime}$ and a randomly sampled value. The rationale behind this practice is due to the spatial correlation found in the prediction from image generative neural networks. This fact makes the direct learning of the important pixel as a blob of connected pixels, which doesn’t distinguish the underlying structure. Comparing the prediction with an uncorrelated random pattern can solve this issue and create isolated pixels that reflect the importance of the underlying structure. The random pattern $P$ is a grayscale image with the same dimensions as C-Image, with its pixel value randomly sampled from a uniform distribution between 0 and 1. We use plastic sampling to generate $P$ for the superior quality of the learned IM compared to other sampling strategies like random sampling, regular sampling, and Halton sampling. The IM can be retrieved through an independent Bernoulli process through rejection sampling.
@@
$IM_{ij}=Sigmoid(\alpha(PIM^{\prime}_{ij}-P_{ij})),\quad Sigmoid(x)=\frac{1}{1+e^{-x}}$ (8)
where $\alpha$ is a positive scalar determining the steepness of the Sigmoid function. The pixel value of each IM is in the range $[0,1]$. The above equation is only used for training the IML Net. When using IML Net during the inferencing, which will be discussed in Section˜2.4.4, only the regular rejection sampling of Eq.˜7 is used. In that case, a binary mask using 1 bit per pixel is enough to represent the importance mask.
@@
During the inferencing stage, this layer is replaced by actual pixel-level rendering computation defined by the chosen volume rendering algorithm for learned important pixels.
@@ 2.3.2 Loss Functions
The IML Net can be trained by itself with a compaction loss function, as shown inside the IML Net block of Fig.˜3, which is an image loss function between the input C-Image and reconstructed C-Image. This practice is suitable for optimizing the off-the-shelf pre-trained RecNNs without going through the long end-to-end training process resulting from adding a complex downstream RecNN. The IML Net can also be trained together with downstream RecNNs selected for various sampling patterns. In this case, the loss function is the weighted sum of the previous compaction loss and the image loss between the predicted reconstructed full-resolution image (FR-Image) and the ground truth (GT) image rendered using the baseline rendering algorithm without using RecNN. This end-to-end training will give better reconstruction results by jointly optimizing the parameters from both the IML Net and the downstream RecNN. In our work, we use the Mean Squared Error (MSE) as the image loss function and 0.5 as the weight for both image losses if trained end-to-end.
@@ 2.4.1 Network Architecture
The focus of this paper is on visualizing scientific data, which is a time-sensitive application. Therefore, we need to keep the main steps of the pipeline, including training and inferencing, finished as fast as possible. Our training includes two steps: training the IML Net and training the IMS Net. We provide a method to speed up the training of IML network, as mentioned in Section˜3.6.4, by training a standalone IML Net without connecting to the downstream RecNN. However, we cannot carry out the same optimization on training the IMS Net. As a result, we have to select a GAN with less training/inferencing complexity. DCGAN [25] provides a better trade-off between the complexity and the reconstruction quality than other GAN architectures like Conditional GAN (cGAN) [14] and StyleGAN [18] whose training/inferencing times are too long for scientific visualization. Fig.˜5 shows the network architecture of IMS Net and the discriminator used for training. IMS Net is a generator or regressor that takes a sparse set of view parameters to generate a 2D binary image as the IM. The IMS Net is trained through the generative adversarial network (GAN). Key components of the network are:
@@
view=(eye_x/y/z,lookAt_x/y/z,up_x/y/z)view=(eye\_x/y/z,\;lookAt\_x/y/z,\;up\_x/y/z) (10)
where the eyeeye is the camera position, $lookAt$ is the focal point where the camera is looking, and $up$ is the "up" direction of the camera. IMS Net is a DCGAN with its detailed structure shown in Fig.˜6(a). It consists of several transposed convolutional layers with batch normalization and ReLU activation functions.
Discriminator: Fig.˜6(b) shows the detailed structure of the binary classifier as a discriminator. The discriminator consisted of several convolutional layers with batch normalization and Leaky ReLU activation functions.
@@ 2.4.2 Loss Functions
As shown in Fig.˜5, we consider two losses to train the IMS Net: image loss on the IM and adversarial loss. The adversarial loss is composed of two parts: the generator loss and the discriminator loss. The discriminator loss $L_{D}$ is the cross-entropy of the binary classifier measuring how well the discriminator can distinguish between real and generated images.
$L_{D}=-\mathbb{E}_{x\sim p_{x}(x)}[\log D(x)]-\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (11)
where $D(\mathbf{x})$ is the discriminator’s output for a real image. $D(G(\mathbf{z}))$ is the discriminator’s output for a fake image $G(\mathbf{z})$. The generator loss $L_{G}$ used is the non-saturating loss measuring how well the generator can fool the discriminator.
@@
$p_{x}(x)$ is the probability distribution of real training data and $p_{z}(z)$ is the probability distribution of the noise vector $z$. The loss function selection and configuration are crucial in training a GAN. Detailed results using different loss functions for training IMS Net are discussed in Section˜3.5.2.
Figure 7: Network architecture of the proposed visualization neural rendering network, IML/S Net + RecNN. IML/S is the trained IMS Net connected with the decoder of the trained IML Net.
@@
Dataset Resolution Size Data Type Chameleon $1024^{2}\times 1088$ 2.1 GB uint16 Beechnut $1024^{2}\times 1546$ 3.0 GB uint16 Rayleigh-Taylor $1024^{3}$ 4.0 GB float32 Flame $1408\times 1080\times 1100$ 6.23 GB float32